Smart Cards have been growing in popularity in recent years. In 2006, 500 million Smart Cards are expected to be issued. This represents a 117% increase from the number of cards that were issued in 2001. How many Smart Cards were issued in 2001? Round to the nearest million.
I set it up as 1.17 · x + x = 500
multiplied all by 100 117 · 200x = 50000
subtracted 117 from both sides 200x = 49883
divided both sides by 200 x = 249.415
The books answer is 230 mil.

Found 2 solutions by joyofmath, Earlsdon:
Answer by joyofmath(189)

Smart Cards have been growing in popularity in recent years. In 2006, 500 million Smart Cards are expected to be issued. This represents a 117% increase from the number of cards that were issued in 2001. How many Smart Cards were issued in 2001? Round to the nearest million.
If the number of cards issued went up by 117% then you have to multiply the 2001 number by 2.17 to get 500,000,000.
So, x*2.17=500000000 where x is the number we're trying to solve for.
Divide both sides by 2.17 and x=230,414,746.54377
That's 230,000,000 when you round to the nearest million.

Answer by Earlsdon(6294)

You have set it up correctly but there is a little problem with the math.
If x represents the number of Smart Cards issued in 2001, then:
x + 1.17x = 500 (in millions) Simplifying this:
2.17x = 500 Multiplying both sides by 100.
217x = 50000 Divide both sides by 217
x = 230.4147...millions. Round to the nearest million.
x = 230 million.